# Leveraging Prototypical Representations for Mitigating Social Bias without Demographic Information

### v1

@@ Abstract
Mitigating social biases typically requires identifying the social groups associated with each data sample. In this paper, we present DAFair, a novel approach to address social bias in language models. Unlike traditional methods that rely on explicit demographic labels, our approach does not require any such information. Instead, we leverage predefined prototypical demographic texts and incorporate a regularization term during the fine-tuning process to mitigate bias in the model’s representations. Our empirical results across two tasks and two models demonstrate the effectiveness of our method compared to previous approaches that do not rely on labeled data. Moreover, with limited demographic-annotated data, our approach outperforms common debiasing approaches.
@@
Shadi Iskander        Kira Radinsky        Yonatan Belinkovthanks:   Supported by the Viterbi Fellowship in the Center for Computer Engineering at the Technion. shadi.isk@campus.technion.ac.il kirar@cs.technion.ac.il        belinkov@technion.ac.il Technion – Israel Institute of Technology
@@ 1 Introduction and Background
Common approaches for social bias mitigation require explicit annotation of biases for each sample in the data Beutel et al. (2017); Zhang et al. (2018). Recent concept removal methods Ravfogel et al. (2020, 2022a, 2022b); Iskander et al. (2023a) have shown promise in addressing social bias by removing sensitive attributes. These approaches rely on training classifiers for predicting the sensitive attribute, and training such classifiers typically requires a significant amount of annotated data.
@@ 2.1.1 Social Attribute Representations
In the absence of labeled data, we leverage semantic similarity and define pairs of texts that capture the models’ understanding of text describing different social attribute groups. For example, to represent gender in an occupation prediction task we can use the encoder’s representations of “This biography is about a man” and “This biography is about a woman”. To generate these pre-defined representations, we employ a generative model. Additional details about the prompt and the resulting pre-defined texts for each task are in Appendix A.1.
@@
 [^1] We call this method SemiDAFair.
When a limited number of labels are available, we leverage the representations generated by the text encoder to derive data-driven representations for each labeled group. Specifically, we calculate the mean representation of each labeled group using the available labeled samples.
@@ 3.3 Metrics
To evaluate extrinsic bias, we align with previous work De-Arteaga et al. (2019); Ravfogel et al. (2020) and use the True Positive Rate Gap (TPR-GAP) as the main fairness metric to assess performance disparities across different protected attribute groups. We also incorporate statistical fairness metrics for a comprehensive evaluation. The metrics along with their calculation procedures is provided in Appendix D.
@@ 3.4 Compared Methods
We compare with two existing methods that do not rely on demographic information: JTT Liu et al. (2021), which trains in a second phase on up-weighed hard examples, and BLIND Orgad and Belinkov (2023), which uses a success detector to down-weigh biased examples.
When only limited labeled samples are available for the sensitive attribute, we adjust and evaluate three methods: INLP Ravfogel et al. (2020), RLACE Ravfogel et al. (2022a), and IGBP Iskander et al. (2023a). These methods operate in a post-hoc manner, aiming to remove demographic information from text representations. More details on these methods can be found in Appendix C.1.2.
@@ 3.5 Settings
In this setting, we explore scenarios where demographic labels are not available. We evaluate the performance of demographic-agnostic methods.
@@
Additionally, we investigate a scenario where we have limited access to demographic labels. We apply various methods while varying the size of the available demographic data to analyze their effectiveness.
DAFair Hyperparameters.
Without demographic labels, validating the selection of prototypical texts and the number of pairs ($K$) is challenging. To mitigate the impact of parameters selection we first generate 10 prototypical texts, then randomly sample $K=4$ pairs within each iteration. We discuss this further in Appendix A.1.
We run each method using 5 random seeds and report the mean and standard deviation of the test results. More details on training and evaluation procedures are described in Appendix C.1.
Occupation Prediction Sentiment Analysis
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Accuracy $\uparrow$ TPR-GAP $\downarrow$
Original
$83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$
JTT
$81.34\pm{\scriptstyle 0.81}$ $14.19\pm{\scriptstyle 1.08}$ $78.08\pm{\scriptstyle 1.21}$ $23.67\pm{\scriptstyle 1.87}$
BLIND
$82.52\pm{\scriptstyle 0.23}$ $13.76\pm{\scriptstyle 1.18}$ $76.45\pm{\scriptstyle 0.67}$ $22.68\pm{\scriptstyle 2.40}$
@@
4 Results and Analysis
4.1 No Demographic Labels
Table 1 presents the evaluation results on the occupation prediction and twitter sentiemnt tasks using BERT as encoder. In both tasks, Our proposed method, DAFair, achieves a slightly lower accuracy compared to the finetuned model. However, it significantly reduces the TPR-GAP, outperforming BLIND and JTT in mitigating bias related to gender or race. This demonstrates the effectiveness of DAFair in addressing bias while maintaining competitive performance in the downstream task. Evaluation with other statistical fairness metrics reveals similar patterns to TPR-GAP (Appendix E). Results with the DeBERTA-V3 model exhibit the same trend, as presented in Appendix E.
@@ 4.2 Limited Demographic Labels
Figure 2 presents TPR-GAP for both tasks under different levels of labeled data for social attributes. While using no labels (horizontal solid lines), DAFair outperforms other methods even when they are provided a limited number of labels (up to 100 in twitter sentiment and 1000 in occupation prediction). DAFair further benefits from labels (SemiDAFair lines), even outperforming prior methods with limited labeled data. With an abundance of labeled data (1000 in sentiment and 100K in occupation prediction), other methods perform better.
@@ 5 Conclusion
We introduced DAFair, a novel approach for mitigating social bias in language models without explicit demographic information. Our method leverages semantic similarity to manipulate the model’s text representations and promote fairness. Experimental results on two tasks demonstrated the effectiveness of DAFair in reducing bias while maintaining competitive performance, even with limited or no labeled demographic data. Our approach offers a practical and flexible solution for bias mitigation in real-world applications, contributing to the development of fairer language models.
@@
Appendix A DAFair Implementation Details
A.1 Pre-defined Representations
To generate pairs of prototypical texts, we provided ChatGPT OpenAI (2022) with a description of the approach, DAFair, along with a description of each dataset and task, and instructed the model to produce 10 pairs of prototypical texts for each dataset. The prototypical texts are presented in Tables 2 and 3.
A.2 DAFair Hyperparameters
Under the setting of no demographic labels, there is no validation set to optimize the selection of protoypical texts or the number of pairs. To avoid dependency on the choice of prototypical representations, we first generate $N>K$ pairs, and within each iteration, we randomly sample $K$ pairs. For all experiments, we set $N=10$ , $K=4$ to capture diverse associations of the training samples with demographic attributes without relying on an extensive set of pairs. In Section F, we analyze the impact of $K$ on the model’s performance and assess its implications on fairness and bias mitigation.
Male Prototypical Texts Female Prototypical Texts
This is a biography about a male.
This is a biography about a female.
A man who excelled in his field.
A woman who excelled in her field.
He is known for his achievements in various industries.
She is known for her achievements in various industries.
A prominent male figure in history.
A prominent female figure in history.
His career and accomplishments are well-regarded.
Her career and accomplishments are well-regarded.
This biography focuses on the life of a distinguished man.
This biography focuses on the life of a distinguished woman.
An influential male individual.
An influential female individual.
He made significant contributions to his profession.
She made significant contributions to her profession.
This is a story about a man who shaped his industry.
This is a story about a woman who shaped her industry.
His impact on his field is noteworthy.
Her impact on her field is noteworthy.
AAE Prototypical Texts SAE Prototypical Texts
This tweet reflects a [sentiment] from a white writer.
This tweet reflects a [sentiment] sentiment from a black writer.
A tweet expressing a [sentiment] moment by a white individual.
A tweet expressing a [sentiment] moment by a black individual.
A [sentiment] viewpoint shared by a writer using Standart American English.
A [sentiment] viewpoint shared by a writer using African American English.
This post, written in standard English, conveys [sentiment] from a white perspective.
This post, written in AAE, conveys [sentiment] from a black perspective.
A message filled with [sentiment] from a white communicator.
A message filled with [sentiment] from a black communicator.
A white person shares their [sentiment] thoughts in this tweet.
A black person shares their [sentiment] thoughts in this tweet.
This is an example of a tweet with [sentiment] in white sociolect.
This is an example of a tweet with [sentiment] sentiment in AAE.
A tweet written by a white speaker that conveys [sentiment].
A tweet written by a black speaker that conveys [sentiment].
This post by a white individual radiates [sentiment] and [sentiment].
This post by a black individual radiates [sentiment] and [sentiment].
A [sentiment] perspective presented by a writer using white sociolect.
A [sentiment] perspective presented by a writer using African American English.
Table 3: Pre-defined Representations for AAE and SAE Tweet Texts
Appendix B Information Removal Methods
INLP Ravfogel et al. (2020) aims to remove linear information from the neural representation by iteratively training a linear classifier to predict the demographic attribute from the neural representation, then projecting the representations to the null space of the linear classifier.
RLACE Ravfogel et al. (2022b) is similar to INLP with the goal of linear information removal from the neural representations of a trained model. However, it uses a different approach based on a linear minimax game.
IGBP Iskander et al. (2023b) overcome the drawbacks of INLP and RLACE which only remove linearly encoded information, and aims to remove non-linear information from representations by gradient-based projections.
Appendix C Setup
C.1 Training Setup
@@
C.1.1 DAFair
To maintain suitable representations in the embedding space, we processed the pre-defined text (Sec A.1) through the text encoder every 200 batches during fine-tuning. This fixed frequency was found to be effective for stable training and did not require further tuning.
@@ C.1.1 DAFair
where $p$ represents a measure of progress, and $\gamma$ controls the rate of change. The parameter $\gamma$ allows us to control the speed at which $\lambda$ approaches $\lambda_{threshold}$. In our experiments, we set $\gamma$ to a fixed value of 5. This schedule allows for stable training in the early stages of training. In order to optimize the performance of DAFair and SemiDAFair, we conducted a grid search to find the optimal value for the parameter $\lambda_{threshold}$. The search was performed over the following values $\lambda_{threshold}\in(0.1,0.2,0.5,1,2,5,10,20,50,100)$.
C.1.2 Compared Methods
@@
C.2 Evaluation Setup
To perform model selection without the need for a validation set with demographic annotations, we adopt Orgad and Belinkov (2023)’s strategy that prioritizes selecting the most radical hyperparameters, while ensuring that the downstream task accuracy remains above 0.97 of the original accuracy.
@@ C.2 Evaluation Setup
Appendix D Fairness Metrics
@@
Appendix E Full Results
Tables 4 and 5 present comprehensive results for the occupation prediction and sentiment analysis tasks, respectively, employing BERT as the text encoder. Each method’s performance is evaluated across multiple metrics, including Accuracy, TPR-GAP, Independence, Separation, and Sufficiency (Section D). Here we also see that our proposed method reduces Independence, Separation, and Sufficiency values, in both tasks.
@@ Appendix E Full Results
The evaluation results for the occupation prediction and sentiment analysis tasks using the DeBERTa-V3 model are presented in Tables 6, and 7. In the two tasks, Both JTT and BLIND methods demonstrate some success in reducing bias, although not substantial. However, DAFair , outperforms both JTT and BLIND in terms of mitigating bias related to social attributes. It achieves a lower TPR-GAP, Independence, Separation and Sufficiency in most cases, while maintaining a comparable level of accuracy. This indicates that our approach is more effective in reducing bias without sacrificing the overall performance of the model also with DeBERTa-V3 model.
@@
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original
$83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $0.16\pm{\scriptstyle 0.002}$ $2.59\pm{\scriptstyle 0.07}$ $2.46\pm{\scriptstyle 0.08}$
JTT
$81.34\pm{\scriptstyle 0.81}$ $14.19\pm{\scriptstyle 1.08}$ $0.16\pm{\scriptstyle 0.002}$ $2.77\pm{\scriptstyle 0.01}$ $2.22\pm{\scriptstyle 0.14}$
BLIND
$82.52\pm{\scriptstyle 0.23}$ $13.76\pm{\scriptstyle 1.18}$ $0.16\pm{\scriptstyle 0.002}$ $2.35\pm{\scriptstyle 0.10}$ $2.21\pm{\scriptstyle 0.09}$
@@
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original
$79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$ $0.17\pm{\scriptstyle 0.003}$ $0.11\pm{\scriptstyle 0.003}$ $0.08\pm{\scriptstyle 0.004}$
JTT
$78.08\pm{\scriptstyle 1.21}$ $23.67\pm{\scriptstyle 1.87}$ $0.18\pm{\scriptstyle 0.01}$ $0.11\pm{\scriptstyle 0.008}$ $0.08\pm{\scriptstyle 0.006}$
BLIND
$76.45\pm{\scriptstyle 0.67}$ $22.68\pm{\scriptstyle 2.40}$ $0.14\pm{\scriptstyle 0.001}$ $0.05\pm{\scriptstyle 0.001}$ $0.06\pm{\scriptstyle 0.001}$
@@
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original
$83.42\pm{\scriptstyle 0.26}$ $14.60\pm{\scriptstyle 0.81}$ $0.17\pm{\scriptstyle 0.003}$ $2.67\pm{\scriptstyle 0.10}$ $2.51\pm{\scriptstyle 0.09}$
JTT
$81.98\pm{\scriptstyle 1.70}$ $14.30\pm{\scriptstyle 0.55}$ $0.18\pm{\scriptstyle 0.004}$ $2.65\pm{\scriptstyle 0.04}$ $2.38\pm{\scriptstyle 0.12}$
BLIND
$82.41\pm{\scriptstyle 0.55}$ $13.86\pm{\scriptstyle 1.49}$ $0.16\pm{\scriptstyle 0.003}$ $2.31\pm{\scriptstyle 0.12}$ $2.29\pm{\scriptstyle 0.10}$
@@
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original
$78.36\pm{\scriptstyle 0.93}$ $29.62\pm{\scriptstyle 1.58}$ $0.19\pm{\scriptstyle 0.003}$ $0.12\pm{\scriptstyle 0.004}$ $0.09\pm{\scriptstyle 0.005}$
JTT
$77.41\pm{\scriptstyle 1.76}$ $28.06\pm{\scriptstyle 2.03}$ $0.18\pm{\scriptstyle 0.013}$ $0.13\pm{\scriptstyle 0.009}$ $0.08\pm{\scriptstyle 0.005}$
BLIND
$77.20\pm{\scriptstyle 1.22}$ $27.32\pm{\scriptstyle 2.84}$ $0.15\pm{\scriptstyle 0.001}$ $0.07\pm{\scriptstyle 0.001}$ $0.08\pm{\scriptstyle 0.001}$

### v2

@@ Abstract
Mitigating social biases typically requires identifying the social groups associated with each data sample. In this paper, we present DaFair, a novel approach to address social bias in language models. Unlike traditional methods that rely on explicit demographic labels, our approach does not require any such information. Instead, we leverage predefined prototypical demographic texts and incorporate a regularization term during the fine-tuning process to mitigate bias in the model’s representations. Our empirical results across two tasks and two models demonstrate the effectiveness of our method compared to previous approaches that do not rely on labeled data. Moreover, with limited demographic-annotated data, our approach outperforms common debiasing approaches. [^1] Our code is available at https://github.com/technion-cs-nlp/DAFair
@@
Shadi Iskander        Kira Radinsky        Yonatan Belinkov shadi.isk@campus.technion.ac.il kirar@cs.technion.ac.il        belinkov@technion.ac.il Technion – Israel Institute of Technology
@@ 1 Introduction and Background
Studies showed that language models have the ability to capture demographic information about the writer, including race or gender, within their representations Caliskan et al. (2017); Zhao et al. (2018). However, this capability can introduce unintended biases, leading to discriminatory outputs De-Arteaga et al. (2019).
@@
Common approaches for social bias mitigation require explicit annotation of biases for each sample in the data Beutel et al. (2017); Zhang et al. (2018). Recent concept removal methods Ravfogel et al. (2020, 2022a, 2022b); Iskander et al. (2023) have shown promise in addressing social bias by removing sensitive attributes. These approaches rely on training classifiers for predicting the sensitive attribute, and training such classifiers typically requires a significant amount of annotated data.
@@ 2.1.1 Social Attribute Representations
In the absence of labeled data, we leverage semantic similarity and define pairs of texts that capture the models’ understanding of text describing different social attribute groups. For example, to represent gender in an occupation prediction task we can use the encoder’s representations of “This biography is about a man” and “This biography is about a woman”. To generate these pre-defined representations, we employ a generative model. We provided ChatGPT OpenAI (2022) with a description of the approach, DaFair, along with a description of each dataset and task, and instructed the model to produce 10 pairs of prototypical texts for each task. The prototypical texts (Tables 7 and 8) and the full prompt (Figure 4) are provided in the appendix.
@@
When a limited number of labels are available, we leverage the representations generated by the text encoder to derive data-driven representations for each labeled group. Specifically, we calculate the mean representation of each labeled group using the available labeled samples. We call this method Semi-DaFair.
@@ 3.3 Metrics
To evaluate extrinsic bias, we align with previous work De-Arteaga et al. (2019); Ravfogel et al. (2020) and use the True Positive Rate Gap (TPR-GAP) as the main fairness metric to assess performance disparities across different protected attribute groups. Following the guidelines in Orgad and Belinkov (2022) for a comprehensive evaluation, we also incorporate statistical fairness metrics: Independence, Separation and Sufficiency. The metrics details and calculation procedures are provided in Appendix B.
Occupation Prediction Sentiment Analysis
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Accuracy $\uparrow$ TPR-GAP $\downarrow$
Original $83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$
JTT $81.34\pm{\scriptstyle 0.81}$ $14.19\pm{\scriptstyle 1.08}$ $78.08\pm{\scriptstyle 1.21}$ $23.67\pm{\scriptstyle 1.87}$
BLIND $82.52\pm{\scriptstyle 0.23}$ $13.76\pm{\scriptstyle 1.18}$ $76.45\pm{\scriptstyle 0.67}$ $22.68\pm{\scriptstyle 2.40}$
@@ 3.4 Compared Methods
We compare with two existing methods that do not rely on demographic information:
JTT Liu et al. (2021), which trains in a second phase on up-weighed hard examples.
BLIND Orgad and Belinkov (2023), which uses a success detector to down-weigh biased examples.
When only limited demographic labeled samples are available, we evaluate three methods:
INLP Ravfogel et al. (2020) removes linear information from the neural representation by iteratively training a linear classifier to predict the demographic attribute from the representation, then projecting the representations to the null-space of the linear classifier.
RLACE Ravfogel et al. (2022b) is similar to INLP with the goal of linear information removal from the neural representations. However, it uses a different approach of a linear minimax game.
IGBP Iskander et al. (2023) overcome the drawbacks of INLP and RLACE which only remove linearly encoded information, and removes non-linear information from representations by gradient-based projections.
@@ 3.5 Settings
In this setting, we explore scenarios where demographic labels are not available. We evaluate the performance of demographic-agnostic methods: JTT, BLIND and DAFAIR.
@@
Additionally, we investigate a scenario where we have limited access to demographic labels. In this setting, we apply information removal methods along with Semi-DaFairwhile varying the size of the available demographic-labeled data to analyze their effectiveness.
We run each method using 5 random seeds and report the mean and standard deviation of the test results. More details on training setup and evaluation procedures are described in Appendix A.
3.6 DaFair Hyperparameters
Under the setting of no demographic labels, there is no validation set to optimize the selection of prototypical texts or the number of pairs. To avoid dependency on the choice of prototypical representations, we first generate $N>K$ pairs, and within each iteration, we randomly sample $K$ pairs. For all experiments, we set $N=10$ , $K=4$ to capture diverse associations of the training samples with demographic attributes, without relying on an extensive set of pairs. In Section 4.3, we analyze the impact of $K$ on the model’s performance and assess its implications on fairness and bias mitigation.
$\lambda$ Tuning.
To perform $\lambda$ tuning without the need for a validation set with demographic annotations, we adopt Orgad and Belinkov (2023)’s strategy that prioritizes selecting the most radical parameter, while ensuring that the downstream task accuracy remains above 0.97 of the original accuracy. More details are described in Appendix A.
4 Results and Analysis
4.1 No Demographic Labels
Table 1 presents the evaluation results on the occupation prediction and twitter sentiemnt tasks using BERT as encoder. In both tasks, Our proposed method, DaFair, achieves a slightly lower accuracy compared to the finetuned model. However, it significantly reduces the TPR-GAP, outperforming BLIND and JTT in mitigating bias related to gender or race. This could be attributed to the fact that JTT and BLIND do not directly address social bias, but assign different weights to examples based on their difficulty. In contrast, DaFair uses a regularization term designed to lower the association of text representation with specific social groups, which might explain the superior reduction of social bias measures. Evaluation with other statistical fairness metrics reveals similar patterns to TPR-GAP (Appendix C). Results with the DeBERTA-V3 model exhibit same trend, as presented in Appendix C.
@@ 4.1 No Demographic Labels
Occupation Prediction Sentiment Analysis
K Accuracy $\uparrow$ TPR-GAP $\downarrow$ Accuracy $\uparrow$ TPR-GAP $\downarrow$
Original $83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$
1 $82.85\pm{\scriptstyle 0.18}$ $13.11\pm{\scriptstyle 0.62}$ $77.52\pm{\scriptstyle 1.23}$ $22.69\pm{\scriptstyle 1.81}$
2 $82.55\pm{\scriptstyle 0.23}$ $12.84\pm{\scriptstyle 0.59}$ $77.16\pm{\scriptstyle 1.12}$ $22.21\pm{\scriptstyle 1.08}$
4 $82.32\pm{\scriptstyle 0.13}$ $12.29\pm{\scriptstyle 0.32}$ $77.20\pm{\scriptstyle 1.17}$ $21.72\pm{\scriptstyle 0.82}$
8 $82.29\pm{\scriptstyle 0.30}$ $12.20\pm{\scriptstyle 0.25}$ $77.47\pm{\scriptstyle 0.94}$ $22.17\pm{\scriptstyle 1.17}$
@@ 4.2 Limited Demographic Labels
Figure 2 presents TPR-GAP for both tasks under different levels of labeled data for social attributes, showcasing the performance of various debiasing methods (Section 3.4), including our proposed methods, DaFair and Semi-DaFair. [^3] Accuracy figures are available in Appendix C.1. While using no labels (horizontal solid lines), DaFair outperforms other methods even when they are provided a limited number of labels (up to 100 in twitter sentiment and 1000 in occupation prediction). DaFair further benefits from labels (Semi-DaFair lines), even outperforming prior methods with limited labeled data. With an abundance of labeled data (1000 in sentiment and 100K in occupation prediction), other methods perform better.
The aim of information removal methods is to learn and neutralize decision boundaries between different social attributes using supervised learning. However, limited labeled examples may hinder classifiers from modeling social attribute subspace in high-dimensional spaces. In Semi-DaFair, we aim to mitigate associations with specific social groups. Surprisingly, a small set of labeled data seems sufficient for this purpose, as more labeled data does not offer additional benefits.
4.3 Effect of Number of Prototypical Texts
To investigate the effect of the number of prototypical text pairs (K) on model performance, we conducted experiments with varying K values of (1, 2, 4, 8). The results presented in Table 2 reveal that all K values contributes to the reduction of the TPR-GAP without affecting accuracy. While larger values of K result in more substantial reductions, the incremental improvements become less significant for $K>2$. These findings suggest that a small K may be sufficient for DaFair.
@@ 5 Conclusion
We introduced DaFair, a novel approach for mitigating social bias in language models without explicit demographic information. Our method leverages semantic similarity to manipulate the model’s text representations during finetuning to promote fairness. Experimental results on two tasks and under different settings demonstrated the effectiveness of DaFair in reducing bias and improving fairness while maintaining competitive downstream task performance, even with limited or no labeled demographic data. With its focus on social bias, DaFair offers a flexible framework adaptable to address other forms of bias through the modification of prototypical texts.
In conclusion, our approach offers a practical and flexible solution for bias mitigation in real-world applications, contributing to the development of fairer language models.
@@
Acknowledgements
This research was supported by the Israel Science Foundation (grant 448/20), an Azrieli Foundation Early Career Faculty Fellowship, and an AI Alignment grant from Open Philanthropy.
Appendix A Setup
A.1 Training Setup
@@
A.1.1 DaFair
To maintain suitable representations in the embedding space, we processed the pre-defined text (Sec 2.1.1) through the text encoder every 200 batches during fine-tuning. This fixed frequency was found to be effective for stable training and did not require further tuning.
@@ A.1.1 DaFair
where $p$ represents a measure of progress, and $\gamma$ controls the rate of change. The parameter $\gamma$ allows us to control the speed at which $\lambda$ approaches $\lambda_{threshold}$.
In our experiments, we set $\gamma$ to a fixed value of 5. This schedule allows for stable training in the early stages of training. In order to optimize the performance of DaFair and Semi-DaFair, we conducted a grid search to find the optimal value for the parameter $\lambda_{threshold}$. The search was performed over the following values $\lambda_{threshold}\in(0.1,0.2,0.5,1,2,5,10,20,50,100)$.
A.1.2 Compared Methods
@@
A.2 Evaluation Setup
As discussed in Section 3.6, to perform parameter tuning without the need for a validation set with demographic annotations, we adopt Orgad and Belinkov (2023)’s strategy that prioritizes selecting the most radical hyperparameters, while ensuring that the downstream task accuracy remains above 0.97 of the original accuracy. This is done for all methods.
@@ A.2 Evaluation Setup
Appendix B Fairness Metrics
@@
Appendix C Full Results
Tables 3 and 4 present comprehensive results for the occupation prediction and sentiment analysis tasks, respectively, employing BERT as the text encoder. Each method’s performance is evaluated across multiple metrics, including Accuracy, TPR-GAP, Independence, Separation, and Sufficiency (Section B). Here we also see that our proposed method reduces Independence, Separation, and Sufficiency values, in both tasks.
@@ Appendix C Full Results
The evaluation results for the occupation prediction and sentiment analysis tasks using the DeBERTa-V3 model are presented in Tables 5, and 6. In the two tasks, Both JTT and BLIND methods demonstrate some success in reducing bias, although not substantial. However, DaFair , outperforms both JTT and BLIND in terms of mitigating bias related to social attributes. It achieves a lower TPR-GAP, Independence, Separation and Sufficiency in most cases, while maintaining a comparable level of accuracy. This indicates that our approach is more effective in reducing bias without sacrificing the overall performance of the model also with DeBERTa-V3 model.
@@
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original $83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $0.16\pm{\scriptstyle 0.002}$ $2.59\pm{\scriptstyle 0.07}$ $2.46\pm{\scriptstyle 0.08}$
JTT $81.34\pm{\scriptstyle 0.81}$ $14.19\pm{\scriptstyle 1.08}$ $0.16\pm{\scriptstyle 0.002}$ $2.77\pm{\scriptstyle 0.01}$ $2.22\pm{\scriptstyle 0.14}$
BLIND $82.52\pm{\scriptstyle 0.23}$ $13.76\pm{\scriptstyle 1.18}$ $0.16\pm{\scriptstyle 0.002}$ $2.35\pm{\scriptstyle 0.10}$ $2.21\pm{\scriptstyle 0.09}$
@@
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original $79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$ $0.17\pm{\scriptstyle 0.003}$ $0.11\pm{\scriptstyle 0.003}$ $0.08\pm{\scriptstyle 0.004}$
JTT $78.08\pm{\scriptstyle 1.21}$ $23.67\pm{\scriptstyle 1.87}$ $0.18\pm{\scriptstyle 0.01}$ $0.11\pm{\scriptstyle 0.008}$ $0.08\pm{\scriptstyle 0.006}$
BLIND $76.45\pm{\scriptstyle 0.67}$ $22.68\pm{\scriptstyle 2.40}$ $0.14\pm{\scriptstyle 0.001}$ $0.05\pm{\scriptstyle 0.001}$ $0.06\pm{\scriptstyle 0.001}$
@@
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original $83.42\pm{\scriptstyle 0.26}$ $14.60\pm{\scriptstyle 0.81}$ $0.17\pm{\scriptstyle 0.003}$ $2.67\pm{\scriptstyle 0.10}$ $2.51\pm{\scriptstyle 0.09}$
JTT $81.98\pm{\scriptstyle 1.70}$ $14.30\pm{\scriptstyle 0.55}$ $0.18\pm{\scriptstyle 0.004}$ $2.65\pm{\scriptstyle 0.04}$ $2.38\pm{\scriptstyle 0.12}$
BLIND $82.41\pm{\scriptstyle 0.55}$ $13.86\pm{\scriptstyle 1.49}$ $0.16\pm{\scriptstyle 0.003}$ $2.31\pm{\scriptstyle 0.12}$ $2.29\pm{\scriptstyle 0.10}$
@@
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original $78.36\pm{\scriptstyle 0.93}$ $29.62\pm{\scriptstyle 1.58}$ $0.19\pm{\scriptstyle 0.003}$ $0.12\pm{\scriptstyle 0.004}$ $0.09\pm{\scriptstyle 0.005}$
JTT $77.41\pm{\scriptstyle 1.76}$ $28.06\pm{\scriptstyle 2.03}$ $0.18\pm{\scriptstyle 0.013}$ $0.13\pm{\scriptstyle 0.009}$ $0.08\pm{\scriptstyle 0.005}$
BLIND $77.20\pm{\scriptstyle 1.22}$ $27.32\pm{\scriptstyle 2.84}$ $0.15\pm{\scriptstyle 0.001}$ $0.07\pm{\scriptstyle 0.001}$ $0.08\pm{\scriptstyle 0.001}$